# LAFA: Agentic LLM-Driven Federated Analytics over Decentralized Data Sources

### v1

@@ III-B Design overview
To address these challenges, LAFA follows a modular design philosophy, consisting of a hierarchical multi-agent framework layered on top of an FA engine. LAFA is structured around three main entities: the queriers, the target clients, and the server, and three core stages: hierarchical decomposition, DAG optimization, and FA execution, as illustrated in Figure 1. Each entity plays a specialized role in the FA process:
@@
Figure 2: The workflow of LAFA
@@ III-D Execution workflow
To illustrate LAFA’s full workflow, consider the example shown in Figure 2 of the paper. Queries are submitted in natural language by end-users, as in the example, a data analyst is curious about the average salary in a university and the difference between that of professors and Ph.D. students. This complex query contains multiple sub-intents that need to be accurately decomposed and jointly answered without duplicating computations. It is received by the server and passed first to the coarse-grained planner. It performs the rough decomposition of the query into several sub-queries. Each sub-query is processed by the fine-grained planner with FA operations DAG knowledge. It will identify the task type of each sub-query and decompose it into a preliminary DAG of corresponding FA operations, which includes preprocessing (data filtering), encryption, aggregation, noise addition, decryption, and postprocessing (result computation).
The output preliminary DAGs for each sub-query are then processed by the DAG optimizer to generate an optimized DAG. Before this optimization, there were repeated operations in each preliminary DAG. For example, in the Figure 2, the salary information of both professors and Ph.D. students is repeatedly collected, encrypted, and aggregated. Repeated execution of FA operations like encrypted aggregation and MPC decryption is extremely resource-intensive. Naïve FA execution may involve many rounds, each costing MBs of data transfer, minutes of client CPU load, and TBs of server load[13]. Efficient query planning and operation fusion are thus critical to avoid this waste. The DAG optimizer rewrites and unites these preliminary DAGs into an optimized DAG without repeated operations. It achieves this by identifying the same FA operations and partitioning the clients into groups based on similar features. For example, in the Figure 2, the people in the university will be separated into professors, Ph.D. students, and others. By collecting information from these partitioned parties only once and final calculations, our system achieves the same goal but with less resource usage.
@@ IV-D Ablation study
The operation count comparison in Figure 3 between the variant without the DAG optimizer and our full system reveals substantial efficiency gains: Access to data, encryption, and aggregation operations per client are all reduced from 3.75 to 2.20, demonstrating the effectiveness of the DAG optimizer in reusing intermediate results across sub-queries. The cost of ensuring DP and performing decryption is also lowered from 5.65 to 3.70, indicating that the DAG optimizer consolidates privacy operations rather than redundantly applying them to each isolated output. With all these resource-intensive operations being reduced, the lightweight calculation operation count increases from 3.65 to 5.50. This demonstrates that LAFA can offload complexity to lightweight post-processing to realize more computation-efficient execution.
@@ V-A LLM agents-assisted data analytics
The emergence of LLMs has significantly advanced the landscape of automated data analytics. Recent developments have extended LLM usage from passive tools into active agent roles that can reason, plan, execute, and interact with data environments across multiple operations. ReAct (Reason + Act) [1] introduces a foundational prompting paradigm where LLMs interleave reasoning traces with executable actions, enabling them to solve complex tasks by interacting with external environments like Wikipedia or code sandboxes. It has inspired many subsequent agent-based designs in data analytics, especially those requiring iterative exploration and interaction.
@@ V-B Federated analytics
FA has become a prominent approach for enabling privacy-preserving data analytics. Existing FA research addresses a range of query types, including those based on count-mean-sketch techniques [10, 19], frequency-related tasks [20, 21, 22], and graph metrics applications qin2016heavy [23, 24, 25, 26].
@@ VI-B Example queries
    What is the average salary of all individuals, what is the difference in average salary between male and female individuals, and how does it compare between those with a bachelor’s degree and those without a college degree?
@@
    What is the average capital gain for all individuals, how does it compare between individuals in executive managerial roles and those in blue-collar jobs, and how does it differ between married and single individuals?
  3. 3.
  4. 4.
@@
  3. 3.
    What is the average acceptance rate of QuickType suggestions across all users, and how does it vary for users typing in more than one language?
  4. 4.
    What are the most common newly typed words added to QuickType predictions this week, and how often are they accepted on first suggestion?

### v2

@@ III-B Design overview
To address these challenges, LAFA follows a modular design philosophy, consisting of a hierarchical multi-agent framework layered on top of an FA engine. LAFA is structured around three main entities: the queriers, the target clients, and the server, and three core stages: hierarchical decomposition, DAG optimization, and FA execution, as illustrated in Fig. 1. Each entity plays a specialized role in the FA process:
@@
Figure 2: The workflow of LAFA using a query as an example
@@ III-D Execution workflow
To illustrate LAFA’s full workflow, consider the example shown in Fig. 2 of the paper. Queries are submitted in natural language by end-users, as in the example, a data analyst is curious about the average salary in a university and the difference between that of professors and Ph.D. students. This complex query contains multiple sub-intents that need to be accurately decomposed and jointly answered without duplicating computations. It is received by the server and passed first to the coarse-grained planner. It performs the rough decomposition of the query into several sub-queries. Each sub-query is processed by the fine-grained planner with FA operations DAG knowledge. It will identify the task type of each sub-query and decompose it into a preliminary DAG of corresponding FA operations, which includes preprocessing (data filtering), encryption, aggregation, noise addition, decryption, and postprocessing (result computation).
The output preliminary DAGs for each sub-query are then processed by the DAG optimizer to generate an optimized DAG. Before this optimization, there were repeated operations in each preliminary DAG. For example, in the Fig. 2, the salary information of both professors and Ph.D. students is repeatedly collected, encrypted, and aggregated. Repeated execution of FA operations like encrypted aggregation and MPC decryption is extremely resource-intensive. Naïve FA execution may involve many rounds, each costing MBs of data transfer, minutes of client CPU load, and TBs of server load[13]. Efficient query planning and operation fusion are thus critical to avoid this waste. The DAG optimizer rewrites and unites these preliminary DAGs into an optimized DAG without repeated operations. It achieves this by identifying the same FA operations and partitioning the clients into groups based on similar features. For example, in the Fig. 2, the people in the university will be separated into professors, Ph.D. students, and others. By collecting information from these partitioned parties only once and final calculations, our system achieves the same goal but with less resource usage.
@@ IV-D Ablation study
The operation count comparison in Fig. 3 between the variant without the DAG optimizer and our full system reveals substantial efficiency gains: Access to data, encryption, and aggregation operations per client are all reduced from 3.75 to 2.20, demonstrating the effectiveness of the DAG optimizer in reusing intermediate results across sub-queries. The cost of ensuring DP and performing decryption is also lowered from 5.65 to 3.70, indicating that the DAG optimizer consolidates privacy operations rather than redundantly applying them to each isolated output. With all these resource-intensive operations being reduced, the lightweight calculation operation count increases from 3.65 to 5.50. This demonstrates that LAFA can offload complexity to lightweight post-processing to realize more computation-efficient execution.
@@ V-A LLM agents-assisted data analytics
The emergence of LLMs has significantly advanced the landscape of automated data analytics. Recent developments have extended LLM usage from passive tools into active agent roles that can reason, plan, execute, and interact with data environments across multiple operations[16]. ReAct (Reason + Act) [1] introduces a foundational prompting paradigm where LLMs interleave reasoning traces with executable actions, enabling them to solve complex tasks by interacting with external environments like Wikipedia or code sandboxes. It has inspired many subsequent agent-based designs in data analytics, especially those requiring iterative exploration and interaction.
@@ V-B Federated analytics
FA has become a prominent approach for enabling privacy-preserving data analytics. Existing FA research addresses a range of query types, including those based on count-mean-sketch techniques [10, 20], frequency-related tasks [21, 22, 23, 24], and graph metrics analytics and learning [25, 26, 27, 28, 29].